OpenStudy (teller):

A line passes through the point −6, 1 and has a slope of 4 over 3. Write an equation in slope-intercept form for this line.

OpenStudy (anonymous):

does "4 over 3" mean \(\frac{4}{3}\)?

OpenStudy (teller):

Yes..

OpenStudy (anonymous):

ok use \[y-y_1=m(x-x_1)\] with \[x_1=-6,y_1=1,m=\frac{4}{3}\]

OpenStudy (teller):

Confused... But I will try it 6-1^1= 4/3(x-x^1) ??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i see you are a bit confused

OpenStudy (misty1212):

you do know that the answer you are supposed to get has an \(x\) and a \(y\) in it right?

OpenStudy (teller):

Yes I'm confused and no .That's why I'm asking for help .

OpenStudy (misty1212):

it should look like \(y=mx+b\) for some \(m\) and \(b\)

OpenStudy (misty1212):

the \(x\) and \(y\) are variables in this expression \[y-y_1=m(x-x_1)\] they are the \(x\) and \(y\) in the answer the \(m\) and \(x_1,y_1\) are numbers @satellite73 showed you what those numbers are

OpenStudy (teller):

So y-1=4/3(x-(-6)

OpenStudy (misty1212):

that looks good

OpenStudy (misty1212):

\[y-1=\frac{4}{3}(x+6)\] looks even better since \(-(-6)=+6\)

OpenStudy (misty1212):

you are not done yet though, you still have to solve for \(y\) which takes exactly two steps

OpenStudy (teller):

So how do I do that?

OpenStudy (misty1212):

a) distribute the \(\frac{4}{3}\) on the right b) add \(1\) to both sides

OpenStudy (misty1212):

does that make any sense to you?

OpenStudy (teller):

y=4/3x+9??

OpenStudy (misty1212):

yes!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (teller):

Mind if you help me with another one?

OpenStudy (misty1212):

sure and say hi to penn for me

OpenStudy (teller):

Hello penn?.?.Don't know who that is lol but Hi... A line passes through the point -2, -4 and has a slope of −7 . Write an equation in slope-intercept form for this line.

OpenStudy (anonymous):

this one is even easier because there is no fraction

OpenStudy (anonymous):

same as before \[y+4=-7(x+2)\]

OpenStudy (misty1212):

what @satellite73 said, then you have to solve for \(y\) using the same two steps a) distribute on the right b) subtract 4 from both sides

OpenStudy (teller):

y=-7x-18??

OpenStudy (misty1212):

https://www.youtube.com/watch?v=oJhYySXzOq0

OpenStudy (misty1212):

yeah looks good

OpenStudy (teller):

A line passes through the point 7, 5 and has a slope of 2 . Write an equation in slope-intercept form for this line. Would it look like this ? y+5=2(x+7)

OpenStudy (misty1212):

not quite

OpenStudy (misty1212):

the reason they were plus before was because each number was negative but the formula is \[y-y_1=m(x-x_1)\]

OpenStudy (misty1212):

so change those plus signs to minus signs

OpenStudy (teller):

so each time it's a plus it will be a minus and when ever it's a minus it will be a plus?

OpenStudy (misty1212):

yes

OpenStudy (teller):

so y-5=2(x+7) or is y-5=2(x-7)

OpenStudy (misty1212):

second one

OpenStudy (teller):

so y=2x-9?

OpenStudy (misty1212):

looks good

OpenStudy (teller):

A line passes through the point −8, 7 and has a slope of 5/4 . Write an equation in slope-intercept form for this line. y+7=5/4(x+8)

OpenStudy (misty1212):

not quite there is one mistake in a sign do you see it?

OpenStudy (teller):

would it be (x-8)??

OpenStudy (misty1212):

no that part is right

OpenStudy (teller):

so -7??

OpenStudy (misty1212):

yes

OpenStudy (teller):

so y-7=5/4(x+8)

OpenStudy (misty1212):

at the risk of repeating myself, it is \[\huge y-y_1=m(x-x_1)\] the minus signs are in the formula

OpenStudy (misty1212):

yeah that is right

OpenStudy (teller):

y=5/4x+17??

OpenStudy (misty1212):

yup

OpenStudy (teller):

it said it was incorrect...

OpenStudy (teller):

it was minus 17..

OpenStudy (misty1212):

no it isn't

OpenStudy (teller):

That's what the problem said and when i hit explain it gave me the answer...

OpenStudy (misty1212):

\[y-7=\frac{5}{4}(x+8)\\ y=7=\frac{5}{4}x+10\\ y=\frac{5}{4}x+17\]

OpenStudy (teller):

sometimes i think the thing is rigged ...

OpenStudy (misty1212):

was the point\((7,-8)\)?

OpenStudy (teller):

no -8,7

OpenStudy (misty1212):

yeah sorry that is what i meant

OpenStudy (misty1212):

if the point is \((-8,7)\) then it should be \(y=\frac{5}{4}x+17\)

OpenStudy (misty1212):

tell them it is a mistake

OpenStudy (teller):

A line passes through the point −6, -1and has a slope of −5/2 . Write an equation in slope-intercept form for this line. Would it be y+1=-5/2(x+6)

OpenStudy (misty1212):

yes

OpenStudy (teller):

so it would be y=-5/2x-16??

OpenStudy (misty1212):

hmm lets check

OpenStudy (misty1212):

\[y+1=-\frac{5}{2}(x+6)\]\[y+1=-\frac{5}{2}x-15\]\[y=-\frac{5}{2}x-16\]

OpenStudy (teller):

A line passes through the point 6, 1 and has a slope of 2. Write an equation in slope-intercept form for this line. so y-1=2(x-6)?? y=2x-11??

OpenStudy (misty1212):

yes

OpenStudy (teller):

One more problem then I will be done with that subject.I greatly appreciate you for helping me ! A line passes through the point 4, -9 and has a slope of −4 .Write an equation in slope-intercept form for this line. y+9=-4(x-4) or is it y+9=-4(x+4)

OpenStudy (misty1212):

fist one

OpenStudy (misty1212):

first

OpenStudy (teller):

so the final answer would be y=-4x+7

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

you done with them all?

OpenStudy (teller):

said I was incorrect...

OpenStudy (misty1212):

ok lets check

OpenStudy (misty1212):

A line passes through the point (4, -9) and has a slope of −4 right?

OpenStudy (misty1212):

\[y+9=-4(x-4)\\ y+9=-4x+16\\ y=-4x+7\]

OpenStudy (misty1212):

it is right, if that was the question

OpenStudy (teller):

We are given that the line passes through the point −4, 9 . So, the equation must be true when =x4 and =y−9 . =−9+−44b We can solve this equation for the y -intercept b . −9 =+−44b −9 =+−16b 7 =b We now write the equation of the line in slope-intercept form. =y+−4x7 Here is the answer. =y+−4x7

OpenStudy (teller):

that's what it said

OpenStudy (teller):

but thank you for the help..

OpenStudy (misty1212):

i have read that twice now and it does not make any sense to me

OpenStudy (misty1212):

what does =y+−4x7 even mean?